OpenStudy (cicilybailey14):

Simplify 9(x - 3)=-2x+39

OpenStudy (anonymous):

1 Expand 9x−27=−2x+39 2 Add 27 to both sides 9x=−2x+39+27 3 Simplify −2x+39+27 to −2x+66 9x=−2x+66 4 Add 2x to both sides 9x+2x=66 5 Simplify 9x+2x to 11x 11x=66 6 Divide both sides by 11 x=6611 7 Simplify 6611 to 6 x=6

OpenStudy (anonymous):

;)

OpenStudy (cicilybailey14):

omg thank you so much lol...I thought U sucked at math??